OpenStudy (anonymous):

2) The population P in 1996 for a state is given along with r, its annual % rate of continuous growth P=8 Millions, r = 2.2% Estimate the population in 2017 = ?

OpenStudy (anonymous):

\[P(t) = Po(1-r)^t \]\[P(t) = 8(1-0.022)^t \]\[P(t) = 8 (0.978)^t \]\[P(2017-1996) = 8(0.978)^{2017-1996}\]

OpenStudy (anonymous):

so the population in 2017 is?

OpenStudy (anonymous):

you tell me ^^

OpenStudy (anonymous):

\[8(0.978)^{21}\]

OpenStudy (anonymous):

oh I see now

OpenStudy (anonymous):

1) In 2000 the population of a state was 4 million and growing continuously at a rate of 3.4% per year. Estimate the state’s population in 2005 = ?

OpenStudy (anonymous):

how would I do this one?

OpenStudy (anonymous):

same way... you have Po = 4 million, r = (3.4)/100 and t = 2005 - 2000

OpenStudy (anonymous):

so I would have 4(3.4/100)^5

OpenStudy (anonymous):

sorry i typed the fomula wrong instead of - it's +\[P(t)=Po(1+r)^t\]

OpenStudy (anonymous):

almost you'd have \[P(t)=4(1+{3.4 \over 100})^5\]

OpenStudy (anonymous):

don't forget to add 1 to the rate r

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

np